0.1x+0.7y=-2.3
0.3x-0.2y=2.3
solve the system using substitution method

Answer by mananth(16949)

0.1x+0.7y=-2.3
0.3x-0.2y=2.3
multiply both equations by 10
x+7y=-23
3x-2y=23
x+7y=-23 .............1
-2y=23-3 x---------2
/-2
y=-11.5- 1.5x
Plug the value of y in (1)
x+7(-11.5+1.5x)= -23
x-80.5+10.5x)=-23
x+10.5x= -23
11.5x=57.5
/11.5
x=5
Plug the value of x in (1)
x+7y=-23
5+7y=-23
5+7y=-23
7y=-28
/7
y=-4